# video-SALMONN 2: Captioning-Enhanced Audio-Visual Large Language Models

### v1

@@ Abstract
Videos contain a wealth of information, and generating detailed and accurate descriptions in natural language is a key aspect of video understanding. In this paper, we present video-SALMONN 2, an advanced audio-visual large language model (LLM) with low-rank adaptation (LoRA) designed for enhanced video (with paired audio) captioning through directed preference optimisation (DPO). We propose new metrics to evaluate the completeness and accuracy of video descriptions, which are optimised using DPO. To further improve training, we propose a novel multi-round DPO (MrDPO) approach, which involves periodically updating the DPO reference model, merging and re-initialising the LoRA module as a proxy for parameter updates after each training round (1,000 steps), and incorporating guidance from ground-truth video captions to stabilise the process. Experimental results show that MrDPO significantly enhances video-SALMONN 2’s captioning accuracy, reducing the captioning error rates by 28%. The final video-SALMONN 2 model, with just 7 billion parameters, surpasses leading models such as GPT-4o and Gemini-1.5-Pro in video captioning tasks, while maintaining highly competitive performance to the state-of-the-art on widely used video question-answering benchmarks among models of similar size. Codes are available at https://github.com/bytedance/video-SALMONN-2.
@@ 1 Introduction
Large language models (LLMs) have exhibited outstanding capabilities in a wide range of natural language processing tasks, and in some instances, have even approached human-level performance [1, 2, 3, 4, 5, 6]. LLMs’ remarkable ability to understand, generate, and reason with text has sparked widespread interest among researchers, attracting both academia and industry to extend them to multimodal understanding and generation. To endow LLMs with multimodal understanding capability, recent studies adopted a paradigm of training modality adapters and aligners between multimodal encoders and LLMs. This approach leverages world knowledge in the textual LLM to interpret diverse types of data perceived by multimodal encoders, enabling the generation of meaningful insights. Over the past two years, many multimodal LLMs have emerged following this paradigm across different modalities. These include models for image and silent video understanding [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], audio understanding [20, 21, 22, 23, 24, 25, 26, 27], and audio-visual understanding [28, 29, 30, 31, 32, 33, 34].
Text descriptions paired with data are essential for building effective multimodal large language models (LLMs). This is because most contemporary multimodal LLMs treat captioning as a cornerstone task during pre-training or supervised fine-tuning (SFT), aligning the output space of multimodal encoders with the input space of textual LLM backbones. Such alignment enables the models to recognise and understand events in multimodal data. Therefore, collecting high-quality, detailed, and low-hallucination text descriptions aligned with multimodal inputs is crucial for enhancing the model’s multimodal understanding and reasoning capabilities. In the context of video understanding, generating accurate and comprehensive captions remains a significant challenge. Videos encode complex content that spans both spatial features within individual frames and audio-visual events across temporal sequences. Despite the prevalence of incompleteness and hallucinations in video captions, few studies have addressed caption quality, largely due to the lack of reliable quantitative evaluation metrics and effective training methods. Moreover, although audio is typically paired with video and provides vital complementary information, most existing visual LLMs lack audio processing capabilities, further contributing to the deficiencies in existing video captioning systems.
In this paper, we introduce video-SALMONN 2, a multimodal LLM that supports both audio and visual inputs and primarily focuses on detailed and holistic audio-visual captioning. Building upon a pre-trained visual LLM, video-SALMONN 2 is further enhanced with auditory capabilities by training on audio-only data as well as videos with synchronised audio tracks. This enables the model to simultaneously “see” and “hear” the video, emulating the way humans perceive and interpret multimedia content. To accurately assess the performance of the model, new metrics to evaluate captioning quality are proposed, which then serve as the objective to optimise during reinforcement learning (RL) based on direct preference optimisation (DPO) [35]. A novel multi-round DPO (MrDPO) method is proposed and performed based on the preferences guided by the metrics. Experiments demonstrate that video-SALMONN 2 with 7 billion (B) parameters can generate detailed and accurate video captions and even outperforms much larger commercial models such as GPT-4o and Gemini-1.5-Pro, and it also maintains competitive performance to the open-source state-of-the-art (SOTA) multimodal LLMs of similar model size on commonly used video question-answering (QA) benchmarks. Our main contributions can be summarised as follows:
  • We develop video-SALMONN 2, a powerful audio-visual LLM that generates high-quality video captions, outperforming larger commercial models such as GPT-4o and Gemini-1.5 in terms of completeness and accuracy.
  • We introduce an evaluation pipeline that computes the missing and hallucination rates of audio-visual events in video captions using text-based LLMs, breaking down the process into sub-tasks suited for current LLMs. Additionally, we provide a new benchmark for video captioning with a human-annotated test set.
@@
  • We propose the MrDPO approach to optimise multimodal LLMs for video captioning, incorporating periodic updates to the DPO reference model, merging and reinitialising the low-rank adaptation (LoRA) [36] proxy, and smoothing the training loss using SFT based on ground-truth captions.
@@ 2.1 Multimodal LLMs
Following the paradigm of connecting multimodal encoders to LLMs using modality adapters, various models have been developed. For image-based LLMs, LLaVA [7, 8] applies instruction tuning [37] to enhance performance on zero-shot tasks. BLIP-2 [9] uses Q-Former to link a frozen encoder with an LLM, while VILA [11] explores pre-training strategies, achieving impressive results in video QA. InternVL [12] scales up the size of visual encoders for improved image representation. For silent video understanding, Video-LLaVA [13] aligns both image and video adapters to learn unified representations. ShareGPT4Video [14] uses GPT-4 to generate dense video captions, improving data quality, and LLaVA-Hound [38] introduces DPO to enhance video LLMs’ understanding capabilities. LLaVA-OneVision [15] enhances the performance of open multimodal LLMs in single-image, multi-image, and silent video scenarios and enables strong transfer learning from images to videos. LLaVA-Video [16] further creates a high-quality synthetic dataset to further push the performance boundary of video understanding. Qwen2-VL [39] and Qwen2.5-VL [17] transform images of varying resolutions into different numbers of tokens and design rotary position embedding to enhance video understanding. NVILA [40] further enhances the performance of video understanding through scaling up the spatial and temporal resolutions first and then compressing. AuroraCap [41] focuses on video captioning from different perspectives and achieves favourable results on its proposed VDC benchmark. F-16 [18] improves the performance of video LLMs through high-frame-rate modeling with dense sampling at 16 frames per second.
@@
In the realm of audio perception, SALMONN [21] uses a dual-encoder structure and can perform zero-shot audio reasoning tasks. LTU [24] and LTU-AS [25] trained on a large audio event QA dataset can answer open-ended questions about audio content. Qwen-Audio [22] and Qwen2-Audio [23] are built on large amounts of audio data to achieve high performance on a wide range of carefully selected audio tasks. Other works [27, 26] extend the LLM to perceive spatial audio information obtained from microphone array recordings.
@@
As the visual frame sequence is often paired with audio in real-world video recordings, some studies investigate understanding non-silent video. Vid2Seq [42] utilises speech transcriptions to enhance video captioning. video-SALMONN [30] uses a multi-resolution causal Q-Former to understand audio and video simultaneously. video-SALMONN-o1 [34] enhances audio-visual reasoning abilities through process DPO. The Google Gemini model achieves video understanding as a native multimodal LLM built upon text, audio, and visual tokens [28]. AVicuna [33] achieves audio-visual temporal understanding by introducing pseudo-untrimmed video data. Video-LLaMA [43] and Video-LLaMA 2 [29] directly concatenate audio and visual tokens for joint audio and video understanding. InternVideo2 [44] aligns video to audio events, speech and text through cross-modal contrastive learning for joint audio-video understanding. Qwen2.5-Omni [45] not only achieves audio-visual understanding but is also able to generate text and speech in a streaming manner.
@@ 2.2 RL for LLMs
RL with human feedback (RLHF) [46] is a widely adopted strategy for improving the performance of textual LLMs. Early approaches commonly employed proximal policy optimisation [47] in conjunction with a reward model trained on human preference data. Building on this foundation, DPO [35] eliminates the need for a separate reward model by leveraging the LLM itself to optimise directly from paired preference data. KTO [48] further simplifies the process by removing the requirement for paired preference data altogether. Extending this direction, RLAIF [49] adopts a cost-efficient framework by using model-generated feedback in place of human input, significantly reducing human involvement. Most recently, GRPO [50] improves efficiency and stability, particularly in tasks such as mathematical reasoning, by comparing relative rewards among candidate responses within a group, thereby removing the need for a separate value network.
@@ 3.1 Model Architecture
Figure 1: The architecture of video-SALMONN 2. The input video is processed by separate visual and audio branches, which extract visual and audio tokens from the corresponding frame sequences. These tokens are then synchronously interleaved and combined with tokens from the text prompt to form the input to the LLM backbone. During training, video-SALMONN 2 dynamically merges and reinitialises LoRA proxies, ensuring that only one LoRA proxy is active at any given time.
In this structure, audio and visual tokens are computed independently in their respective branches. For the visual branch, the input visual frame sequence is first downsampled at a fixed frame rate of $\phi$ frames/second, and the total number of frames to sample is $n=\phi\,T$, where $T$ is the duration of the input video in seconds. Let $m$ be the maximum number of frames to sample based on the resource constraint. If $n>m$, the frame rate is further reduced to $\phi^{\prime}=\lfloor m/T\rfloor$, resulting in $n=\phi^{\prime}T\leq m$. Let $\mathbf{I}_{i}$ be the $i$th sampled visual frame, and each visual frame in $\mathbf{I}_{1},\mathbf{I}_{2},\ldots,\mathbf{I}_{n}$ is transformed to visual tokens independently using a pre-trained visual encoder $\text{Encoder}_{\text{Visual}}$ followed by a visual modality aligner $\text{Aligner}_{\text{Visual}}$, as shown below:
$\mathbf{H}^{\text{Visual}}_{i}=\text{Aligner}_{\text{Visual}}(\text{Encoder}_{\text{Visual}}(\mathbf{I}_{i})),~{}~{}1\leq i\leq n,$
where $\mathbf{H}^{\text{Visual}}_{i}$ is the visual tokens corresponding to $\mathbf{I}_{i}$.
The audio frame sequence $\mathbf{S}$ is fed into a pre-trained audio encoder $\text{Encoder}_{\text{Audio}}$. Since $\text{Encoder}_{\text{Audio}}$ may have a maximum processing duration $t_{\text{max}}$, the audio will be sliced into $l=\lceil{T}/{t_{\text{max}}\rceil}$ segments of $t_{\text{max}}$-length and processed separately by $\text{Encoder}_{\text{Audio}}$ by
$\mathbf{Z}^{\text{Audio}}_{j}=\text{Encoder}_{\text{Audio}}(\mathbf{S}_{(j-1)\times t_{\text{max}}:j\times t_{\text{max}}}),~{}~{}1\leq j\leq l,$
where $\mathbf{Z}^{\text{Audio}}_{j}$ is the audio feature vector output by the audio encoder of the $j$th audio segment.
As suggested by [51], a segment-level positional embedding is added before the modality aligner to improve the performance of long-form audio. Denote $\mathbf{Z}^{\text{Pos}}_{j}$ as the segment-level position embedding matrix corresponding to the position $j$, $\text{Concat}(\cdot)$ as the concatenation operation along the time dimension, and $\text{Aligner}_{\text{Audio}}$ as the audio modality aligner. The audio token sequence $\mathbf{H}^{\text{Audio}}$ for the whole audio can be computed as shown in Eqns. (1)–(3):
$\tilde{\mathbf{Z}}^{\text{Audio}}_{j}$ $=\mathbf{Z}^{\text{Audio}}_{j}+\mathbf{Z}^{\text{Pos}}_{j},~{}~{}1\leq j\leq l$ (1)
$\tilde{\mathbf{Z}}^{\text{Audio}}$ $=\text{Concat}(\tilde{\mathbf{Z}}^{\text{Audio}}_{1},\tilde{\mathbf{Z}}^{\text{Audio}}_{2},\ldots,\tilde{\mathbf{Z}}^{\text{Audio}}_{l})$ (2)
$\mathbf{H}^{\text{Audio}}$ $=\text{Aligner}_{\text{Audio}}(\tilde{\mathbf{Z}}^{\text{Audio}}).$ (3)
Next, the audio and visual tokens are interleaved chronologically to form the input audio-visual token sequence $\mathbf{H}$ fed into the LLM backbone, and $\mathbf{H}$ is obtained based on Eqns. (4)–(6) by
$\alpha_{i}$ $=l\cdot{i}/{n},~{}~{}1\leq i\leq n$ (4)
$\mathbf{H}_{i}$ $=\text{Concat}(\mathbf{H}^{\text{Visual}}_{i},\mathbf{H}^{\text{Audio}}_{\alpha_{i-1}:\alpha_{i}}),~{}~{}1\leq i\leq n$ (5)
$\mathbf{H}$ $=\text{Concat}(\mathbf{H}_{1},\mathbf{H}_{2},\ldots,\mathbf{H}_{n}).$ (6)
Finally, the text-based backbone LLM is required to generate a text response $\mathbf{\hat{Y}}$ given the user’s text prompt $\mathbf{P}$ and the audio-visual token sequence $\mathbf{H}$:
@@
3.2 Training Strategies
To equip the visual LLM with audio perceptual capabilities, we adopt a multi-stage training strategy that enables effective utilisation of audio information for video understanding while preserving the model’s visual processing performance. Starting from a well-trained visual LLM, the training pipeline consists of three stages: audio modality alignment, audio-visual SFT, and RL via the proposed MrDPO method. The backbone LLM, the visual encoder, and the audio encoder are kept frozen during training to prevent catastrophic forgetting.

### v3

# video-SALMONN 2: Caption-Enhanced Audio-Visual Large Language Models

@@ Abstract
We present video-SALMONN 2, a family of audio-visual large language models that set new state-of-the-art (SOTA) results in video description and question answering (QA). Our core contribution is multi-round direct preference optimisation (MrDPO), paired with a caption-quality objective that jointly rewards completeness and factual accuracy. Unlike standard DPO with a fixed reference policy, MrDPO periodically refreshes the reference by bootstrapping from a newly re-initialised lightweight adapter trained on the latest preferences, avoiding reference staleness and enabling continual improvement. This strategy produces captions that are consistently more detailed and accurate than those from proprietary systems such as GPT-4o and Gemini-1.5 Pro. We further distil these gains by using our model to generate a high-quality video-caption corpus for supervised fine-tuning of new models, transferring benefits beyond captioning to strong performance on complex video-QA tasks. Across widely used audio-visual and visual-only understanding benchmarks (including Video-MME, WorldSense, AVUT, Video-Holmes, DailyOmni, MLVU, and LVBench), our 3B and 7B models achieve SOTA results at comparable scales, while the 72B model surpasses all other open-source systems. Our source code, models, and data are released at https://github.com/bytedance/video-SALMONN-2.
@@ 1 Introduction
The remarkable success of large language lodels (LLMs) in text-based tasks, where they have approached human-level performance, has inspired widespread efforts to extend their capabilities to multimodal understanding (OpenAI et al., 2024; Dubey et al., 2024; Yang et al., 2025a). A prevalent strategy is to train modality adapters/aligners that bridge pretrained multimodal encoders and a textual LLM, allowing the LLM’s knowledge to interpret non-text inputs and produce meaningful outputs. Over the past two years, numerous multimodal LLMs following this paradigm have appeared across modalities, including image and silent video (Liu et al., 2024b; a; Li et al., 2023; Bai et al., 2023; Lin et al., 2023; Chen et al., 2023; Lin et al., 2024; Chen et al., 2024; Li et al., 2024; Zhang et al., 2024b; Bai et al., 2025; Li et al., 2025; Zhang et al., 2025), audio (Wu et al., 2023; Tang et al., 2024b; Chu et al., 2023; 2024; Gong et al., 2024; 2023; Tang et al., 2024c; Zheng et al., 2024), and audio–visual understanding (Team et al., 2024; Cheng et al., 2024; Sun et al., 2024; Fu et al., 2024; 2025b; Tang et al., 2024d; Sun et al., 2025).
High-quality text descriptions paired with data are fundamental to training effective multimodal LLMs. Contemporary systems often use captioning as a core objective in pre-training or supervised fine-tuning (SFT), aligning multimodal encoder outputs with a textual LLM’s input space and thereby enabling event recognition and reasoning. Thus, collecting detailed, faithful, low-hallucination captions that are tightly aligned with the inputs is critical for robust multimodal understanding. Yet video captioning remains difficult: videos couple rich intra-frame spatial content with audio–visual events evolving over time. Progress has been limited by unreliable quantitative metrics and limited training strategies for caption quality and by the common practice of discarding the audio stream despite its complementary value. These gaps not only impair caption generation but also constrain downstream performance on general tasks such as video question answering (QA).
We introduce video-SALMONN 2, a family of multimodal LLMs that ingest both audio and visual inputs with a primary focus on detailed, holistic audio–visual captioning. Building on a pretrained visual LLM, video-SALMONN 2 is trained on audio-only corpora and on videos with paired audio tracks, enabling the model to see and hear simultaneously and to capture fine-grained temporal interactions across modalities. To accurately assess and optimise caption quality, we propose new captioning metrics that evaluate completeness and factuality and use them as reward signals for reinforcement learning (RL) via direct preference optimisation (DPO) (Rafailov et al., 2024). We further introduce multi-round DPO (MrDPO), which iteratively updates the reference policy using a lightweight adapter, driving continual improvements in captioning. Finally, we use the MrDPO-trained model to generate higher-quality captions and perform SFT on this improved data to train new models. Experiments show that our final models retain state-of-the-art (SOTA) captioning performance and transfer gains to general video understanding. On widely used audio–visual and visual-only video QA benchmarks, including Video-MME, WorldSense, AVUT, Video-Holmes, DailyOmni, MLVU, and LVBench, our final models with 3 billion (B) and 7B parameters achieve SOTA results at comparable scales, and our 72B model outperforms proprietary systems such as GPT-4o and Gemini-1.5 Pro. Our main contributions can be summarised as follows:
  • We develop video-SALMONN 2, a family of strong audio–visual LLMs with enhanced video-captioning ability that surpasses proprietary systems such as GPT-4o and Gemini-1.5 Pro.
  • We introduce a lightweight evaluation pipeline that uses text-only LLMs to estimate missing and hallucinated audio–visual events by decomposing caption assessment into LLM-friendly sub-steps, enabling metric-driven RL for caption optimisation. We also release a human-annotated video-captioning benchmark to validate our evaluation approach.
@@
  • We propose MrDPO, a caption-optimisation method that periodically updates the DPO reference policy by merging and re-initialising a low-rank adaptation (LoRA) (Hu et al., 2022) proxy, and stabilises training by smoothing the loss with SFT on ground-truth captions.
  • We show that gains from video captioning can transfer to video QA. An MrDPO-enhanced captioner yields a higher-quality SFT corpus, and models trained on it (video-SALMONN-2+) achieve new open-source SOTA results across multiple video-QA benchmarks.
@@ 2.1 Multimodal LLMs
Following the adapter-based paradigm that connects multimodal encoders to LLMs, a broad family of models has emerged. For images, LLaVA applies instruction tuning to boost zero-shot performance (Liu et al., 2024b; a; Wei et al., 2022); BLIP-2 links a frozen vision encoder to an LLM via the Q-Former (Li et al., 2023); VILA studies pre-training strategies with strong video-QA transfer (Lin et al., 2023); and InternVL scales visual encoders for stronger image representations (Chen et al., 2023). For silent video, Video-LLaVA aligns image- and video-specific adapters to learn unified representations (Lin et al., 2024); ShareGPT4Video uses GPT-4 to generate dense captions to improve data quality (Chen et al., 2024); and LLaVA-Hound introduces DPO to enhance video understanding (Zhang et al., 2024a). More recent systems expand capability and scale: LLaVA-OneVision strengthens open-source multimodal LLMs across single-/multi-image and silent-video settings with strong image-to-video transfer (Li et al., 2024); LLaVA-Video constructs a high-quality synthetic corpus to push video understanding further (Zhang et al., 2024b); Qwen2-VL and Qwen2.5-VL map variable-resolution images to variable token counts and adopt rotary position embeddings for improved video modelling (Wang et al., 2024a; Bai et al., 2025); NVILA scales spatial/temporal resolution before compression (Liu et al., 2024c); AuroraCap targets video captioning from multiple perspectives and reports favourable results on the proposed VDC benchmark (Chai et al., 2025); and F-16 improves performance via high-frame-rate dense sampling at 16 frame per second (FPS) (Li et al., 2025).
@@
In the realm of audio perception, SALMONN (Tang et al., 2024b) uses a dual-encoder structure and can perform zero-shot audio reasoning tasks. LTU (Gong et al., 2024) and LTU-AS (Gong et al., 2023) trained on a large audio event QA dataset can answer open-ended questions about audio content. Qwen-Audio (Chu et al., 2023) and Qwen2-Audio (Chu et al., 2024) are built on large amounts of audio data to achieve high performance on a wide range of carefully selected audio tasks. Other works (Zheng et al., 2024; Tang et al., 2024c) extend the LLM to perceive spatial audio information obtained from microphone array recordings.
@@
As the visual frame sequence is often paired with audio in real-world video recordings, some studies investigate understanding non-silent video. Vid2Seq (Yang et al., 2023) utilises speech transcriptions to enhance video captioning. video-SALMONN (Sun et al., 2024) uses a multi-resolution causal Q-Former to understand audio and video simultaneously. video-SALMONN-o1 (Sun et al., 2025) enhances audio-visual reasoning abilities through process DPO. The Google Gemini model achieves video understanding as a native multimodal LLM built upon text, audio, and visual tokens (Team et al., 2024). AVicuna (Tang et al., 2024d) achieves audio-visual temporal understanding by introducing pseudo-untrimmed video data. Video-LLaMA (Zhang et al., 2023) and Video-LLaMA 2 (Cheng et al., 2024) directly concatenate audio and visual tokens for joint audio and video understanding. InternVideo2 (Wang et al., 2024b) aligns video to audio events, speech, and text through cross-modal contrastive learning for joint audio-video understanding. Qwen2.5-Omni (Xu et al., 2025a) not only achieves audio-visual understanding but is also able to generate text and speech in a streaming manner. Qwen3-Omni (Xu et al., 2025b) further enhances omni understanding by supporting longer audio input, enabling multimodal thinking mode, and significantly reducing streaming latency.
@@ 2.2 RL for LLMs
RL with human feedback (RLHF) (Ouyang et al., 2022) is a widely adopted strategy for improving the performance of textual LLMs. Early approaches commonly employed proximal policy optimisation (Schulman et al., 2017) in conjunction with a reward model trained on human preference data. Building on this foundation, DPO (Rafailov et al., 2024) eliminates the need for a separate reward model by leveraging the LLM itself to optimise directly from paired preference data. Iterative RPO (Pang et al., 2024) combines DPO and NLL loss and iteratively conducts optimisation for better reasoning capability. KTO (Ethayarajh et al., 2024) further simplifies the process by removing the requirement for paired preference data altogether. Extending this direction, RL with AI feedback (Lee et al., 2023) adopts a cost-efficient framework by using model-generated feedback in place of human input, significantly reducing human involvement. Most recently, GRPO (Shao et al., 2024) improves efficiency and stability, particularly in tasks such as mathematical reasoning, by comparing relative rewards among candidate responses within a group, thereby removing the need for a separate value network.
@@ 3.1 Model Architecture
Figure 1: The architecture of video-SALMONN 2. The input video is processed by separate visual and audio branches, which extract visual and audio tokens from the corresponding frame sequences. These tokens are then synchronously interleaved and combined with tokens from the text prompt to form the input to the LLM backbone. In the MrDPO stage, a LoRA proxy is re-initialised, trained, and merged into the LLM backbone at the end of each training round.
Specifically, the model accepts a synchronised audio and video stream as its primary input. The video input is first sampled into discrete frames. Each frame is processed independently by a visual encoder to extract salient feature representations. Following this, a visual aligner module maps these high-level visual features into the LLM’s input space, transforming them into a sequence of visual tokens. In a parallel process, the accompanying audio waveform is fed into an audio encoder to capture its corresponding features. Similar to the visual branch, an audio aligner then projects these audio features into the same LLM input space, producing a sequence of audio tokens. Capitalising on the inherent temporal alignment of audio and video signals, the audio tokens and visual tokens are arranged in an interleaved manner. In particular, tokens corresponding to the same one-second segment of the input are grouped together, ensuring that the visual and auditory information for any given moment are adjacent in the sequence. The LLM’s position embeddings are then applied to this combined sequence, creating the final audio-visual token stream that is fed into the model.
In addition to the audio-visual stream, the model takes a user-provided text prompt as input. This prompt is converted into text tokens using the LLM’s text embeddings. These text tokens are then combined with the interleaved audio-visual token sequence, forming the complete input for the LLM. Finally, the text-based backbone LLM is required to generate a text response $\mathbf{\hat{Y}}$ given the user’s text prompt $\mathbf{P}$ and the audio-visual token sequence $\mathbf{H}$:
@@
3.2 Multi-Stage Training for Audio–Visual Captioning
To equip the visual LLM with strong video captioning capabilities, we adopt a multi-stage training strategy that enables effective utilisation of audio information for video understanding while preserving the model’s visual processing performance. Starting from a well-trained visual LLM, the training pipeline mainly consists of three stages: audio modality alignment, audio-visual SFT, and RL via the proposed MrDPO method. The backbone LLM, the visual encoder, and the audio encoder are kept frozen during training to prevent catastrophic forgetting.